(a) Evaluate∫[

x

√x

2+1

]dx.


(b) Use MATLAB to generate some typical integral curves of f(x) =

x

√x

2+1

over the


interval (−5,5).


1
Expert's answer
2021-12-13T14:31:28-0500

a)

"\int\frac{x}{\sqrt{x^2+1}}dx=\int\frac{d(x^2)}{2\sqrt{x^2+1}}dx=\sqrt{x^2+1}+c"


b)

syms x
f=sqrt(x^2+1);
int(f, x, -5, 5)